d^2y/dx^2+4dy/dx-4y=sin3x




1
Expert's answer
2022-01-11T13:19:44-0500

"y''+4y'-4y=\sin3x"

solution find in form

"y=y_1+y_2"

then

"1) y''+4y'-4y=0\\
\lambda^2+4\lambda-4=0\\
\lambda_1=-2-2\sqrt{2}, \lambda_2=-2+2\sqrt{2}\\
y_1=c_1e^{(-2-2\sqrt{2})x}+c_2e^{(-2+2\sqrt{2})x}\\
2) y_2=a\cos3x+b\sin3x\\
y_2'=-3a\sin3x+3b\cos3x\\
y_2''=-9a\cos3x-9b\sin3x\\"

Plug in the equation

"-9a\cos3x-9b\sin3x-12a\sin3x+\\
+12b\cos3x-4a\cos3x-4b\sin3x=\sin3x\\
\cos3x: -9a+12b-4a=0\\
\sin3x: -9b-12a-4b=1\\
a=\frac{12b}{13}\\
-13b-12a=1, -13b-\frac{144b}{13}=1, -313b=13, \\
b=-\frac{13}{313}, a=-\frac{12}{313}\\
y_2=-\frac{12}{313}\cos3x-\frac{13}{313}\sin3x\\"

Answer:

"y=c_1e^{(-2-2\sqrt{2})x}+c_2e^{(-2+2\sqrt{2})x}-\\
-\frac{12}{313}\cos3x-\frac{13}{313}\sin3x"